# On Erlang mixture approximations for differential equations

### v2

@@ 6.2 Identify kernel parameters
For a given order, $M\in\mathbb{N}_{\geq 0}$, and $t_{h}\in\mathbb{R}_{\geq 0}$, the objective is to determine the values of the coefficients, $c_{m}\in[0,1]$ for $m=0,\ldots,M$, and the rate parameter, $a\in\mathbb{R}_{>0}$, that minimize the integral of the squared error,
@@
while satisfying the constraint that the coefficients must sum to one. We approximate this integral by a left rectangle rule (such that the objective function includes $\alpha(0)$) with $N$ rectangles, and we determine the coefficients and rate parameters by solving the inequality-constrained nonlinear program (NLP)
(120a) $\min_{\{c_{m}\}_{m=0}^{M},a}\quad$ $\phi=\frac{1}{2}\sum_{k=0}^{N-1}(\alpha(t_{k})-\hat{\alpha}(t_{k}))^{2}\Delta t,$
(120b) subject to $\sum_{m=0}^{M}c_{m}=1,$
(120c) $0\leq c_{m}\leq 1,\quad m=0,\ldots,M,$
(120d) $a_{\min}\leq a,$
where the quadrature points are $t_{k}=k\Delta t$ for $k=0,\ldots,N-1$, $\Delta t\in\mathbb{R}_{>0}$ is the width of the rectangles, and $a_{\min}\in\mathbb{R}_{>0}$ is an arbitrary lower bound, which ensures that only positive values of $a$ are feasible. Furthermore, the objective function involves the regular kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, and the $M$’th order Erlang mixture kernel, $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with coefficients $\{c_{m}\}_{m=0}^{M}$ and rate parameter $a$, evaluated at the quadrature points.
6.2.1 Derivatives
Many algorithms for solving NLPs in the form (120) can benefit from analytical expressions for the first- and second-order derivative of the objective function. The derivatives of the objective function are
(121a) $\frac{\partial{\phi}}{\partial{v}}$ $=-\sum_{k=0}^{N-1}(\alpha(t_{k})-\hat{\alpha}(t_{k}))\frac{\partial{\hat{\alpha}}}{\partial{v}}(t_{k})\Delta t,$
(121b) $\frac{\partial^{2}{\phi}}{\partial{v}\partial{w}}$ $=-\sum_{k=0}^{N-1}\left((\alpha(t_{k})-\hat{\alpha}(t_{k}))\frac{\partial^{2}{\hat{\alpha}}}{\partial{v}\partial{w}}(t_{k})-\frac{\partial{\hat{\alpha}}}{\partial{w}}(t_{k})\frac{\partial{\hat{\alpha}}}{\partial{v}}(t_{k})\right)\Delta t,$
where $v$ and $w$ represent either $a$ or $c_{m}$. Furthermore, the first-order derivatives of the Erlang mixture kernel are
(122a) $\frac{\partial{\hat{\alpha}}}{\partial{a}}(t)$ $=\sum_{m=0}^{M}c_{m}\frac{\partial{\ell_{m}}}{\partial{a}}(t),$
(122b) $\frac{\partial{\hat{\alpha}}}{\partial{c_{k}}}(t)$ $=\ell_{k}(t),$ $k$ $=0,\ldots,M,$
and the second-order derivatives are
(123a) $\frac{\partial^{2}{\hat{\alpha}}}{\partial^{2}{a}}(t)$ $=\sum_{m=0}^{M}c_{m}\frac{\partial^{2}{\ell_{m}}}{\partial^{2}{a}}(t),$
(123b) $\frac{\partial^{2}{\hat{\alpha}}}{\partial{a}\partial{c_{k}}}(t)$ $=\frac{\partial{\ell_{k}}}{\partial{a}}(t),$ $k$ $=0,\ldots,M,$
(123c) $\frac{\partial^{2}{\hat{\alpha}}}{\partial{c_{k}}\partial{c_{n}}}(t)$ $=0,$ $k$ $=0,\ldots,M,$ $n$ $=0,\ldots,M.$
Finally, the first- and second-order derivatives of the Erlang kernel with respect to the rate parameter are
(124a) $\frac{\partial{\ell_{m}}}{\partial{a}}(t)$ $=\left(\frac{m+1}{a}-t\right)\ell_{m}(t),$ $m$ $=0,\ldots,M,$
(124b) $\frac{\partial^{2}{\ell_{m}}}{\partial^{2}{a}}(t)$ $=\left(\frac{m+1}{a}-t\right)\frac{\partial{\ell_{m}}}{\partial{a}}(t)-\frac{m+1}{a^{2}}\ell_{m}(t),$ $m$ $=0,\ldots,M.$
@@ 6.3 Theoretical Erlang mixture approximation
We will compare the accuracy obtained with the least-squares approach described in Section 6.2 to that of a reference approach based on the theoretical values of the coefficients from Theorem 4.26. First, we determine the size of the approximation interval, $t_{h}$, such that $1-\beta(t_{h})\approx\epsilon$, e.g., using bisection as described in Section 6.1. Next, for a given order, $M$, we determine the rate parameter by $a=t_{h}/(M+1)$, and we use the theoretical expression (17) to compute the coefficients, $c_{m}$, for $m=0,\ldots,M$. In general, $\epsilon$ should be chosen to be small. Consequently, the coefficients will almost sum to one, and it is not necessary to scale them.
@@ 6.4 Implementation
We implement the evaluation of the objective function in the NLP (120) and its first- and second-order derivatives in C, and we use Matlab’s MEX interface to call the C routines from Matlab. Furthermore, we use Matlab’s fmincon to approximate the solution to the NLP, and we provide analytical first- and second-order derivatives. We exploit that the recursion $\ell_{m}(t)=(at/m)\ell_{m-1}(t)$ for $m=1,\ldots,M$ where $\ell_{0}(t)=ae^{-at}$ is less prone to rounding errors than a straightforward evaluation of the expression (11) in the definition, and we use either double or long double precision when evaluating the Erlang kernels, the Erlang mixture kernels, and their derivatives. We simulate the approximate system of ODEs (73) using either Matlab’s ode45 or ode15s, and depending on the order of the Erlang mixture kernel, we implement the matrices $A$, $B$, and $C$ as either dense or sparse. Furthermore, we provide the analytical Jacobian of the right-hand side function of the approximate system of ODEs (73) to the simulator. We implement the numerical method for non-stiff DDEs described in Appendix B.1 in both Matlab and C (with a MEX interface). The numerical method from Appendix B.2 is implemented in Matlab, we use fsolve to solve the residual equations, and we provide the analytical Jacobian. In all three implementations, the right-hand side function, $f$, the delay function, $h$, and the kernel, $\alpha$, are evaluated in Matlab. Finally, in Section 8, we carry out parallel simulations and kernel approximations using a high-performance computing cluster [14].
@@ 7 Numerical test of the accuracy
In this section, we construct a DDE in the form (2b)–(3) with a known solution and use it to investigate the accuracy of the Erlang mixture approximation and the numerical solution of the approximate system of ODEs (73). Furthermore, we assess the convergence rate of the two numerical methods for solving initial value problems in the form (2)–(3) described in Appendix B. Finally, we compare the accuracy of the Erlang mixture approximation with that of one of the exponential mixture approximations described in [18] for a system of DDEs involving a gamma kernel, which describes chemotherapy-induced myelosuppression [33].
@@ 7.1 Modified logistic equation
(129) $E_{x}$ $=\sum_{n=0}^{N_{x}-1}(\hat{x}(t_{n+1})-x^{*}(t_{n+1}))^{2}\Delta t,$ $E_{\alpha}$ $=\sum_{k=0}^{N_{\alpha}-1}(\hat{\alpha}(t_{k})-\alpha(t_{k}))^{2}\Delta t.$
@@
Here, $\hat{x}$ and $\hat{\alpha}$ denote the approximations. Note that the kernel error, $E_{\alpha}$, is analogous to the objective function in (120a) but may contain a different number of terms. We have used the model parameter values $\sigma=4$ mo-1 and $\kappa=1$ and the dilation parameter $\gamma=10$ mo, and the initial and final times are $t_{0}=0$ mo and $t_{f}=24$ mo. When we determine the Erlang mixture approximation of a given order, $M$, we use $N=10^{3}$ points in the objective function in (120a), and we use $a_{\min}=1$ mo-1. Furthermore, we use a threshold of $\epsilon=10^{-14}$ and a convergence tolerance of $10^{-15}$ to determine the approximation interval length, $t_{h}$. Finally, we solve the NLP (120) using Matlab’s fmincon with an optimality tolerance of $10^{-10}$ and a step size tolerance below machine precision, and we use double precision in the involved kernel evaluations. We also use the identified value of $t_{h}$ in the reference approach described in Section 6.3. We use Matlab’s ode45 with absolute and relative tolerances of $10^{-12}$ to simulate the approximate system of ODEs (73), and we use $N_{x}=24\cdot 10^{3}$ quadrature points to evaluate the state error in (129), which corresponds to a time step size of $\Delta t=10^{-3}$ mo. Furthermore, we use $N_{\alpha}=10^{6}$ quadrature points to evaluate the kernel error. For the results obtained with the numerical methods described in Appendix B, we use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and the above time step size, and for the method for stiff DDEs, we use a function tolerance of $10^{-12}$ and an optimality tolerance of $10^{-8}$ when we solve the involved residual equations with Matlab’s fsolve.
@@
The results demonstrate that for this example, the Erlang mixture approximation obtained with the proposed least-squares approach is several orders of magnitude more accurate than the reference approach based on the theoretical expressions for the coefficients. In both cases, there is a high correlation between the error of the kernel and the error of the approximate state trajectory. For the least-squares approach, the convergence rates are faster than polynomial, and for the reference approach, the convergence rates are quadratic for sufficiently large values of $M$ (see also Figure 9, and note that the errors are squared in (129)). The kernel and state error is higher for $M$ equal to 64 and 128 than for $32$ because the objective function in the NLP (120) uses a coarser resolution in the rectangle rule than the error in (129). Finally, the state error decreases quadratically with the time step size for the two numerical methods described in Appendix B because the involved integral and differential equation have been discretized with first-order methods.
@@
Figure 3: Convergence analysis for the modified logistic equation. The state (top left) and kernel (bottom left) errors for Erlang mixture approximations of different orders, $M$, obtained with the proposed least-squares approach and the reference approach based on the theoretical expressions for the coefficients. The bottom right plot shows the state and kernel errors against each other, and the top right plot shows the state error obtained with the numerical approaches for non-stiff and stiff DDEs for different time step sizes, $\Delta t$. The black solid lines in the left column are proportional to $1/(M+1)^{2}$, and the one in the top right is proportional to $\Delta t^{2}$.
@@ 7.2 Myelosuppression
In this section, we compare the accuracy of the Erlang mixture approximation proposed in this work with an exponential mixture approximation (see Section F) used by Guglielmi and Hairer [18] for a system of DDEs describing myelosuppression. The system describes the evolution of precursor cells, $y:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, which ultimately form granulocytes, $w:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$:
@@
where $\Gamma:\mathbb{C}\rightarrow\mathbb{C}$ is the gamma function. The initial and final times are $t_{0}=0$ h and $t_{f}=700$ h, and the initial states are $y(t)=w_{0}$ for $t\in(-\infty,t_{0}]$, $w(t_{0})=w_{0}$, and $C(t_{0})=A_{0}/V$ where $A_{0}=127$ mg/kg and $V=1.03$ L/kg. We use a threshold of $\epsilon=10^{-3}$ in the computation of $t_{h}$, and we use $N=10^{4}$ points in the objective function. The initial guess of the rate parameter, $a$, is $\kappa$, and we use $a_{\min}=10^{-3}$ h-1. As the system is stiff, we use ode15s to simulate the approximate system of ODEs (73), and we compare to the error of the exponential mixture approximation described in [18], which depends on the tolerance $\epsilon_{r}\in\mathbb{R}_{>0}$ that also indirectly determines the number of terms in the approximation. Lower tolerances lead to more terms, and we consider $\epsilon_{r}\in\{10^{-1},10^{-3},10^{-5},10^{-7},10^{-9},10^{-11}\}$. Furthermore, we compare the state obtained with the Erlang mixture approach to that obtained with $\epsilon_{r}=10^{-11}$ using $N_{x}=700\cdot 10^{3}$ points in time, corresponding to a time step size of $\Delta t=10^{-3}$ h. The remaining parameters are the same as in Section 7.1.
@@
The results are shown in Figure 4, and as for the example in Section 7.1, the least-squares approach described in Section 6 is significantly more accurate than the reference approach based on the theoretical expressions for the coefficients. However, the error is generally higher than for the modified logistic equation (see Figure 3), the convergence rates are no longer faster than polynomial for the least-squares approach, and they are no longer quadratic for the reference approach based on the theoretical expressions for the coefficients. The figure also shows that Guglielmi and Hairer’s approach is significantly more accurate (note the difference in the vertical axes). The error of the Erlang mixture approximation is primarily high for $t$ close to zero. The reason for this may be that the derivative of the true kernel has a singularity at zero whereas for finite values of $M$ and $a$, the derivative of the Erlang mixture approximation does not. This may also describe why the convergence rates of the Erlang mixture approximation are reduced. However, we emphasize that the Erlang mixture approximation can be used for a broader class of kernels (bounded and continuous) than the exponential mixture approximations.
@@
Figure 4: Convergence analysis for the myelosuppression model. The state (top left) and kernel (bottom left) errors for Erlang mixture approximations of different orders, $M$, obtained with the least-squares approach and the reference approach based on theoretical values of the coefficients. The kernel error is shown for Guglielmi and Hairer’s approach in the bottom right, and the true and approximate kernels are shown in the top right for $M=128$ and $\epsilon_{r}=10^{-11}$ (corresponding to 309 terms). The black solid line in the top left is proportional to $1/(M+1)^{2}$ whereas the black dashed line in the bottom left is proportional to $1/(M+1)^{3/2}$.

### v3

@@ 5.2 Identify kernel parameters
For a given order, $M\in\mathbb{N}_{\geq 0}$, and $t_{h}\in\mathbb{R}_{\geq 0}$, we choose the rate parameter $a\in\mathbb{R}_{>0}$ as
(70) $a$ $=\frac{M+1}{t_{h}}.$
Next, the objective is to determine the values of the coefficients, $c_{m}\in\mathbb{R}$ for $m=0,\ldots,M$, that minimize the integral of the squared error over the interval $0,t_{h}]$,
@@
while satisfying the constraint that the coefficients must sum to the integral of $\alpha$ over the same interval. We approximate this integral by a left rectangle rule (such that the objective function includes $\alpha(0)$) with $N$ rectangles, and we determine the coefficients by solving the regularized and constrained quadratic program (QP)
(72a) $\min_{\{c_{m}\}_{m=0}^{M}}\hskip 16.0pt$ $\phi=\frac{1}{2}\sum_{k=0}^{N-1}(\alpha(t_{k})-\hat{\alpha}(t_{k}))^{2}\Delta t+\frac{w}{2}\sum_{m=0}^{M}c_{m}^{2},$
(72b) subject to $\sum_{m=0}^{M}c_{m}=\beta(t_{h}),$
where the quadrature points are $t_{k}=k\Delta t$ for $k=0,\ldots,N-1$, $\Delta t\in\mathbb{R}_{>0}$ is the width of the rectangles, and $w\in\mathbb{R}_{>0}$ is a regularization weight. Furthermore, the objective function involves the regular kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, and the $M$’th order Erlang mixture kernel, $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, with coefficients $\{c_{m}\}_{m=0}^{M}$ and rate parameter $a$, evaluated at the quadrature points.
Proposition 5.3.
The unique solution $c\in\mathbb{R}^{M+1}$ to the regularized and constrained least-squares problem (72) is a column vector given by
(73) $c$ $=(L^{T}L+wI)^{-1}(e\lambda+L^{T}g),$
where $\lambda\in\mathbb{R}$ is a Lagrange multiplier, $e\in\mathbb{R}^{M+1}$ is a column vector of ones, and $L\in\mathbb{R}_{\geq 0}^{N\times M+1}$ and $g\in\mathbb{R}^{N}$ are a matrix and vector whose elements are equal to the basis functions and the kernel $\alpha$ evaluated in the quadrature points, respectively:
(74) $\lambda$ $=\frac{\beta(t_{h})-e^{T}(L^{T}L+wI)^{-1}L^{T}g}{e^{T}(L^{T}L+wI)^{-1}e},$ $L_{km}$ $=\ell_{m}(t_{k}),$ $g_{k}$ $=\alpha(t_{k}),$
for $m=0,\ldots,M$ and $k=0,\ldots,N-1$.
Proof 5.4.
The QP (72) is in the form
(75a) $\min_{c}\hskip 18.0pt\phi$ $=\frac{1}{2}c^{T}(L^{T}L+wI)c+g^{T}Lc,$
(75b) $\mathrm{subject~to}\quad$ $e^{T}c=\beta(t_{h}),$
where $I\in\mathbb{R}^{M+1\times M+1}$ is an identity matrix. This is an equality-constrained quadratic program and the first-order necessary conditions for a solution is that there must exist a Lagrange multiplier, $\lambda$, such that the following linear system of equations is satisfied [38, Sec. 16.1]:
(76) $\begin{bmatrix}L^{T}L+wI&-e\\ -e^{T}\end{bmatrix}\begin{bmatrix}c\\ \lambda\end{bmatrix}$ $=\begin{bmatrix}L^{T}g\\ -\beta(t_{h})\end{bmatrix}$.
The expression for $c$ in (73) is the solution to the first row block. Next, we substitute it into the second row block:
(77) $-e^{T}c$ $=-e^{T}(L^{T}L+wI)^{-1}(e\lambda+L^{T}g)=-\beta(t_{h}).$
The expression for $\lambda$ in (74) is obtained by isolating $\lambda$ in the rightmost equation. Finally, the solution is locally unique because the Hessian $L^{T}L+wI$ is positive definite when $w>0$, i.e., $x^{T}(L^{T}L+wI)x=\|Lx\|_{2}^{2}+w\|x\|_{2}^{2}>0$ for all $x\neq 0$ [38, Thm. 12.6].
@@ 5.3 Theoretical Erlang mixture approximation
We will compare the accuracy obtained with the least-squares approach described in Section 5.2 to that of a reference approach based on the theoretical values of the coefficients from Proposition 4.2. First, we determine the size of the approximation interval, $t_{h}$, such that $\rho(\infty)-\rho(t_{h})\approx\epsilon$, e.g., using bisection as described in Section 5.1. Next, for a given order, $M$, we determine the rate parameter by (70), and we use the theoretical expression in (15) to compute the coefficients, $c_{m}$, for $m=0,\ldots,M$.
@@ 5.4 Implementation
We exploit that the recursion $\ell_{m}(t)=(at/m)\ell_{m-1}(t)$ for $m=1,\ldots,M$ where $\ell_{0}(t)=ae^{-at}$ is less prone to rounding errors than a straightforward evaluation of the expression (13) in the definition. We simulate the approximate system of ODEs (19) using either Matlab’s ode45 or ode15s, and depending on the order of the Erlang mixture kernel, we implement the matrices $A$, $B$, and $C$ as either dense or sparse. Furthermore, we provide the analytical Jacobian of the right-hand side function of the approximate system of ODEs (19) to the simulator. We implement the numerical method for non-stiff DDEs described in Appendix G.1 in both Matlab and C (with a MEX interface). The numerical method from Appendix G.2 is implemented in Matlab, we use fsolve to solve the residual equations, and we provide the analytical Jacobian. In all three implementations, the right-hand side function, $f$, the delay function, $h$, and the kernel, $\alpha$, are evaluated in Matlab. Finally, in Section 7, we carry out parallel simulations and kernel approximations using a high-performance computing cluster [14].
@@ 6 Numerical test of the accuracy
In this section, we construct a DDE in the form (3b)–(4) with a known solution and use it to investigate the accuracy of the Erlang mixture approximation and the numerical solution of the approximate system of ODEs (19). Furthermore, we assess the convergence rate of the two numerical methods for solving IVPs in the form (3)–(4) described in Appendix G. Finally, we compare the accuracy of the Erlang mixture approximation with that of one of the exponential mixture approximations described in [18] for a system of DDEs involving a gamma kernel, which describes chemotherapy-induced myelosuppression [32].
@@ 6.1 Modified logistic equation
(82) $E_{x}$ $=\left(\sum_{n=0}^{N_{x}-1}(\hat{x}(t_{n+1})-x^{*}(t_{n+1}))^{2}\Delta t_{x}\right)^{\frac{1}{2}},$ $E_{\alpha}$ $=\left(\sum_{k=0}^{N_{\alpha}-1}(\hat{\alpha}(t_{k})-\alpha(t_{k}))^{2}\Delta t_{\alpha}\right)^{\frac{1}{2}}.$
@@
Here, $\hat{x}$ and $\hat{\alpha}$ denote the approximations, and $\Delta t_{x}=(t_{f}-t_{0})/N_{x}$ and $\Delta t_{\alpha}=t_{h}/N_{\alpha}$. Note that the kernel error, $E_{\alpha}$, is analogous to the square root of the first term in the objective function in (72a) but may contain a different number of terms. We have used the model parameter values $\sigma=4$ mo-1 and $\kappa=1$ and the dilation parameter $\gamma=10$ mo, and the initial and final times are $t_{0}=0$ mo and $t_{f}=24$ mo. When we determine the Erlang mixture approximation of a given order, $M$, we use $N=10^{3}$ points in the objective function in (72a). Furthermore, we use a threshold of $\epsilon=10^{-14}$ and a convergence tolerance of $10^{-15}$ to determine the approximation interval length, $t_{h}$ (see Table 1). We also use the identified value of $t_{h}$ in the reference approach described in Section 5.3. We use Matlab’s ode45 with absolute and relative tolerances of $10^{-12}$ to simulate the approximate system of ODEs (19), and we use $N_{x}=24\cdot 10^{3}$ quadrature points to evaluate the state error in (82), which corresponds to a time step size of $\Delta t_{x}=10^{-3}$ mo. Furthermore, we use $N_{\alpha}=10^{6}$ quadrature points to evaluate the kernel error. For the results obtained with the numerical methods described in Appendix G, we use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and the above time step size, and for the method for stiff DDEs, we use a function tolerance of $10^{-12}$ and an optimality tolerance of $10^{-8}$ when we solve the involved residual equations with Matlab’s fsolve.
@@
The results demonstrate that for this example, the Erlang mixture approximation obtained with the proposed least-squares approach is several orders of magnitude more accurate than the reference approach based on the theoretical expressions for the coefficients. In both cases, there is a high correlation between the error of the kernel and the error of the approximate state trajectory. For the least-squares approach, the convergence rates are faster than polynomial, and for the reference approach, the convergence rates are linear for sufficiently large values of $M$ (see also Figure SM1). The kernel error is higher for $M$ equal to 128 than for $64$ because the objective function in the QP (72) uses a coarser resolution in the rectangle rule than the error in (82). Finally, the state error decreases linearly with the time step size for the two numerical methods described in Appendix G because the involved integral and differential equation have been discretized with first-order methods.
@@
Figure 2: Convergence analysis for the modified logistic equation. The state (top left) and kernel (bottom left) errors for Erlang mixture approximations of different orders, $M$, obtained with the proposed least-squares approach and the reference approach based on the theoretical expressions for the coefficients. The bottom right plot shows the state and kernel errors against each other, and the top right plot shows the state error obtained with the numerical approaches for non-stiff and stiff DDEs for different time step sizes, $\Delta t$. The black solid lines in the left column are proportional to $1/(M+1)$, and the one in the top right is proportional to $\Delta t$.
Table 1: The Erlang mixture approximation order, $M$, the approximation horizon, $t_{h}$, identified with the bisection method from Section 5.1, and the value of the rate parameter, $a$, obtained with (70) for the problems considered in Section 6 (the first two) and Section 7. For the first two problems, $t_{h}$ is fixed, but $M$ varies. Therefore, $a$ varies as well, and this is marked with a v. The third and fourth problems are the bifurcation analyses with respect to the logistic growth rate, $\sigma$, and the kernel parameter, $\mu_{2}$, respectively, and in the latter case, the shown values correspond to the six values of $\mu_{2}$ considered in Figure 4 and 5. The fifth problem involves the six kernels shown in Figure 7.
Problem $M$ $t_{h}$ $a$
Modified logistic equation (forced) v 4.32 mo v
Myelosuppression v 489.08 h v
Modified logistic equation ($\sigma$) 100 1.36 mo 74.41 mo-1
Modified logistic equation ($\mu_{2}$) 350 0.92 mo 380.75 mo-1
1.22 mo 287.08 mo-1
1.45 mo 242.85 mo-1
1.62 mo 217.31 mo-1
1.91 mo 184.13 mo-1
2.32 mo 151.27 mo-1
Nuclear fission 500 12.13 s 41.32 s-1
12.10 s 41.43 s-1
11.97 s 41.86 s-1
11.69 s 42.87 s-1
10.53 s 47.57 s-1
8.28 s 60.50 s-1
@@ 6.2 Myelosuppression
In this section, we compare the accuracy of the Erlang mixture approximation proposed in this work with an exponential mixture approximation (see Section SM4) used by Guglielmi and Hairer [18] for a system of DDEs describing myelosuppression. The system describes the evolution of precursor cells, $y:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, which ultimately form granulocytes, $w:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$:
@@
where $\Gamma:\mathbb{C}\rightarrow\mathbb{C}$ is the gamma function. The initial and final times are $t_{0}=0$ h and $t_{f}=700$ h, and the initial states are $y(t)=w_{0}$ for $t\in(-\infty,t_{0}]$, $w(t_{0})=w_{0}$, and $C(t_{0})=A_{0}/V$ where $A_{0}=127$ mg/kg and $V=1.03$ L/kg. We use a threshold of $\epsilon=10^{-3}$ in the computation of $t_{h}$ (see Table 1), and we use $N=10^{4}$ points in the objective function. The initial guess of the rate parameter, $a$, is $\kappa$. As the system is stiff, we use ode15s to simulate the approximate system of ODEs (19), and we compare to the error of the exponential mixture approximation described in [18], which depends on the tolerance $\epsilon_{r}\in\mathbb{R}_{>0}$ that also indirectly determines the number of terms in the approximation. Lower tolerances lead to more terms, and we consider $\epsilon_{r}\in\{10^{-1},10^{-3},10^{-5},10^{-7},10^{-9},10^{-11}\}$. Furthermore, we compare the state obtained with the Erlang mixture approach to that obtained with $\epsilon_{r}=10^{-11}$ using $N_{x}=700\cdot 10^{3}$ points in time, corresponding to a time step size of $\Delta t_{x}=10^{-3}$ h. The remaining parameters are the same as in Section 6.1.
@@
The results are shown in Figure 3, and as for the example in Section 6.1, the least-squares approach described in Section 5 is significantly more accurate than the reference approach based on the theoretical expressions for the coefficients. However, the error is generally higher than for the modified logistic equation (see Figure 2), the convergence rates are no longer faster than polynomial for the least-squares approach, and they are no longer linear for the reference approach based on the theoretical expressions for the coefficients. The figure also shows that Guglielmi and Hairer’s approach does have converge rates that are faster than polynomial (note the difference in the vertical axes). The error of the Erlang mixture approximation is primarily high for $t$ close to zero. The reason for this may be that the derivative of the true kernel has a singularity at zero whereas for finite values of $M$ and $a$, the derivative of the Erlang mixture approximation does not. This may also describe why the convergence rates of the Erlang mixture approximation are reduced. However, we emphasize that the Erlang mixture approximation can be used for a broader class of kernels (bounded and continuous) than the exponential mixture approximations. Furthermore, the state error is generally high because of the order of magnitude of the state variables (10103030 for $y$ and $w$) and the length of the simulation interval. A constant error of 1 would correspond to a state error of $E_{x}=\sqrt{700}\approx 26.5$. Even for $M+1=4$, the state error is below this value.
@@
Figure 3: Convergence analysis for the myelosuppression model. The maximum state error (top left) and the kernel error (bottom left) for Erlang mixture approximations of different orders, $M$, obtained with the least-squares approach and the reference approach based on theoretical values of the coefficients. The kernel error is shown for Guglielmi and Hairer’s approach in the bottom right, and the true and approximate kernels are shown in the top right for $M=128$ and $\epsilon_{r}=10^{-11}$ (corresponding to 309 terms). The black solid line in the top left is proportional to $1/(M+1)$ whereas the black dashed line in the bottom left is proportional to $1/(M+1)^{3/4}$.